SpaceRef

SpaceRef


NASA ISS On-Orbit Status 26 June 2013

Status Report From: NASA HQ
Posted: Wednesday, June 26, 2013

image International Space Station (ISS) Management Center (IMC) Daily Summary Report
Increment 36 Operations - Stage 36-6
06/26/13 4:00 pm Central Daylight Time (CDT)
177/21:00 Greenwich Mean Time (GMT)

New Status: Ocular Health: Flight Engineer (FE)-5 Parmitano and FE-6 Nyberg collected blood pressure and vital sign data in support of the Ocular Health payload. They then performed tonometry measurements with each acting as a Crew Medical Officer (CMO) and subject. FE-3 Cassidy was the CMO for both Parmitano and Nyberg to assist in the use of the Fundoscope data collection. The Ocular Health protocol aims to systematically gather physiological data to characterize the Risk of Microgravity-Induced Visual Impairment/Intracranial Pressure on crewmembers assigned to a 6 month ISS increment. Researchers believe that by systematically measuring visual, vascular and central nervous system changes over the course of this experiment, a database could be developed to inform future in-flight research protocols, identify mechanisms, develop countermeasures, and direct clinical monitoring and clinical practice guidelines.

US Extravehicular Activity (EVA) Preparations: Cassidy and Nyberg completed Liquid Cooling and Ventilation Garment (LCVG) fill and EVA Mobility Unit (EMU) loop scrub as well as EMU checkout activities for both EMUs that will be used for the upcoming US EVAs scheduled for July 9th and 16th.

Automated Transfer Vehicle (ATV)-4 Cargo Operations: Ground cargo specialists uplinked next week's unpack message following yesterday's completion of unpack activities for this week. Today the crew used next week's message to resume unloading cargo from and loading trash into ATV4, The crew performed 3 hours of cargo operations and are 55% complete.

Mobile Servicing System (MSS) Status: Last night, Robotics Ground Controllers powered up the MSS and maneuvered the Space Station Remote Manipulator System (SSRMS) to grapple the Special Purpose Dexterous Manipulator System (SPDM). They released the SPDM from Mobile Base System (MBS) Power Data Grapple Fixture (PDGF) 2 and maneuvered SPDM Arm 2 to pick up Robot Micro Conical Tool (RMCT) 2 from the SPDM Tool Holder Assembly (THA) with On-Orbit Replaceable Unit (ORU) Tool Changeout Mechanism (OTCM)-2. Finally the Robotics Ground Controllers maneuvered SPDM Arm 2 away from the THA leaving the SSRMS and SPDM in a park position.

When setting up for release of the SPDM from MBS PDGF2, the Robotics Ground Controllers found that the SPDM Latching End Effector (LEE) appeared to be calibrated but the SPDM LEE load cell offset value was not correct. They corrected this and the SPDM was released nominally, putting it and the SSRMS in a good configuration for today's SPDA Door Checkout.

Today's Planned Tasks
- Biochemical Urine Test [Complete]
- Biochemical Urine Test [Complete]
- URYSIS Hardware Stow [Complete]
- Ocular Health (OH) - Blood Pressure Operations [Complete]
- Daily Planning Conference (S-band) [Complete]
- Post-EVA debrief. Tagup with specialists (S-band) [Complete]
- HMS Tonometry Test Setup [Complete]
- T2 Exercise Video Equipment Setup [Complete]
- Filling (separation) of EDV (KOV) for Elektron [Complete]
- Tonometry Exam [Complete]
- HMS - Tonometry Test - CMO [Complete]
- HMS - Tonometry Test - Subject [Complete]
- Tonometry Exam [Complete]
- СОЖ Maintenance
- Tonometry Stow [Complete]
- Closeout operations with replaceable fluid flow regulator РРЖ panel after EVA [Complete]
- JEMRMS RLT Laptop Activation [Complete]
- ТКГ 418 and DC1 Interface Leak Check [Complete]
- JEMRMS Console Checkout [Complete]
- Photo/TV Camcorder Setup Verification [Complete]
- CIR - Hardware Teardown [Complete]
- JEMRMS RLT Laptop Deactivation
- СО-СУ Transfer Hatch Opening [Complete]
- ATV Cargo Operations [Complete]
- Atmosphere Control and Supply (ACS) High Pressure Oxygen Bypass Configuration [Complete]
- On MCC Go ТГК-СУ Hatch Opening [Complete]
- Initiate EMU Cooling Loop Scrub Part 1 [Complete]
- Initiate EMU Cooling Loop Scrub Part 1 [Complete]
- EVA tool stow, update IMS / r/g 2734, 2735 [Complete]
- ТКГ 418 (DC1) deactivation, air duct installation [Complete]
- Replacement of SM ПФ1-4 Dust Filter Cartridge (ФГБ1ПГО_2_226_1, bag No. 10 (00060210R) [Complete]
- Photo/TV Camcorder Setup Verification [Complete]
- On MCC Go ISS Repress from ТКГ 418 СрПК Section 1 (DC1) (start) [Complete]
- T2 Exercise Video Equipment Stow [Complete]
- ISS O2 Repress from ТКГ 418 (DC2) СРПК Section 1 (terminate) [Complete]
- Post-EVA Orlan-MK Spacesuit Equipment Stowage [Complete]
- ATV Cargo Operations [Complete]
- ИП-1 Flow Sensor Position Check [Complete]
- Initiate EMU Cooling Loop Scrub Part 2 [Complete]
- IMS Update
- ATV Transfer Tagup [Complete]
- ТКГ 418 Docking Mechanism Removal [Complete]
- VZAIMODEISTVIYE. Experiment Ops [Complete]
- ТГК 418 (DC1) Docking Mechanism Disassembly, Installation of 2 hatch tools on the external side of the hatch [Complete]
- Waste and Hygiene Compartment (WHS) Pre-Treat Tank and Pre-Treat Tank Hose R&R
- Oxygen Supply System Reconfig to Initial [Complete]
- Tool Handover to US crew after RS EVA [Complete]
- БНП removal from SM [РО] and stowage (БНП No. 3 п.1 (001473R) → МИМ2_ Пт_302) [Complete]
- US EVA Tool Stow [Complete]
- Set up the first Orlan battery pack 825М3 for discharge and start discharging the first pack (No. 1268861161 (00060664R) [Complete]
- ISS - HAM Radio Session [Complete]
- Urine Transfer from ЕДВ-У (3 ea.) to ТГК 418 (DC1) Rodnik БВ2. ТКГ 418 (DC1) Rodnik БВ1 Tank Connector Flush [Complete]
- Daily Planning Conference (S-band) [Complete]
- Ocular Health (OH) Fundoscope Prep [Complete]
- Ocular Health (OH) Fundoscope Exam [Complete]
- Ocular Health (OH) Fundoscope post-Exam Stow [Complete]

Task List
- CEO Ops [Complete]
- WHC KTO replace [Complete]
- WHC EDV R&R
- Roscosmos Report
- Uragan Photo
- Econ Photo
- EDV Separation
- IFM - FFTD find [In work]

Ground
- Nominal Commanding

Three-Day Look Ahead:
Thursday, 6/27: Ocular Health activities, ATV cargo transfer, Robonaut
Friday 6/28: Robonaut, Spinal Scan, ATV cargo transfer
Saturday, 6/29: Crew off duty

QUICK ISS Status - Environmental Control Group:
Component - Status
Elektron - On
Vozdukh - On
[СКВ] 1 - SM Air Conditioner System ("SKV1") - On
[СКВ] 2 - SM Air Conditioner System ("SKV2") - Off
Carbon Dioxide Removal Assembly (CDRA) Lab - Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 - Operate
Major Constituent Analyzer (MCA) Lab - Off
Major Constituent Analyzer (MCA) Node 3 - Operate
Oxygen Generation Assembly (OGA) - Process
Urine Processing Assembly (UPA) - Standby
Trace Contaminant Control System (TCCS) Lab - Off
Trace Contaminant Control System (TCCS) Node 3 - Full up

Meeting Schedule: Meeting Day Date Houston Time (CDT)
IMMT Thursday June 27 8:00 am
IMMT Monday July 1 CANCELLED
IMMT Tuesday July 2 8:00 am
IMMT Thursday July 4 CANCELLED

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.